OpenStudy (dtan5457):

A 940 ML sample of gas is collected over water at 40.0 C. The total pressure is 85.592 KP. Calculate the volume of the dry gas at 25.0 C and 72.661 KP

OpenStudy (dtan5457):

@aaronq

OpenStudy (dtan5457):

I got the separate dry gas pressure as 77.761 KP, and used the combined gas law (converting all celcius to kelvins )

OpenStudy (dtan5457):

then i solved for volume and I got 1001.444316...but the answer is 963.3 ml, which is still pretty far off.

OpenStudy (dtan5457):

for the volume at 25 degrees celcius and 72.661 KP, i got the dry gas from 72.661-3.169

OpenStudy (dtan5457):

maybe i wasn't suppose to do that?

OpenStudy (dtan5457):

@Callisto

OpenStudy (aaronq):

I'm getting the same thing \(\sf \Large \dfrac{[P^{40^o}_{T}-P^{40^o}_{water}]*V_1}{T_1}=\dfrac{[P^{25^o}_{T}-P^{25^o}_{water}]*V_2}{T_2 }\)

OpenStudy (aaronq):

V_2=1007.23 mL

OpenStudy (dtan5457):

Maybe when they say "Calculate the volume of the dry gas at 25.0 C and 72.661 KP", the dry gas pressure is really at 72.661?

OpenStudy (anonymous):

But what is the answer?

OpenStudy (anonymous):

@dtan5457?

OpenStudy (dtan5457):

963.3 ml

OpenStudy (anonymous):

Ok.